OpenStudy (anonymous):

Show the steps for the solutions! 1. Find x. lg 2x=3 2. Find f'(x) when f(x)=1/(1-3x)Show the steps for the solutions! 1. Find x. lg 2x=3 2. Find f'(x) when f(x)=1/(1-3x)@Mathematics

OpenStudy (agreene):

\[\ln 2x = 3\] Is this the first one?

OpenStudy (anonymous):

That is the first problem, yes..

OpenStudy (anonymous):

wait, lg means 10 is the base D:

OpenStudy (alfie):

As long as I am concerned, "lg" stands for base 10 \[\large \begin{array}{l} {\log _{(10)}}2x = 3\\ 2x = {10^3}\\ x = \frac{{1000}}{2}\\ x = 500 \end{array}\]

OpenStudy (agreene):

For the second one... you can just use the chain rule. \[f'(x) = \frac{3}{(1-3x)^2}\]

OpenStudy (anonymous):

Apparently they want to use the definition of derivate to solve the second problem. I tried, but got the wrong answer. Can someone show how?

OpenStudy (agreene):

So, they want you to use that stupid limit thing?

OpenStudy (anonymous):

I think so :p

OpenStudy (anonymous):

but if you were to use the chain rule, how would that be? I'm not taught in english, so I'm not sure what the chain rule is..

OpenStudy (agreene):

Lame... Remember that it is: \[f'(x)=\lim_{h\rightarrow0}\frac{f(x+h)-f(x)}{h}\] \[\large\lim_{h\rightarrow0}\frac{\frac{1}{1-3x+h}-\frac{1}{1-3x}}{h}\] \[\lim_{h\rightarrow0}\frac{1}{-1-h+6 x+3 h x-9 x^2}\] \[\frac{1}{\lim_{h\rightarrow0}(-1-h+6 x+3 h x-9 x^2)}\] \[-\frac{1}{(1-3x)^2}\] Gah... I messed this up somewhere... it's close but not right.

OpenStudy (agreene):

I hate these stupid limits, there are much easier ways of doing them: The Chain Rule says we can do it this way: \[f'(x)=-\frac{\frac{d}{dx}(3-3x)}{(1-3x)^2}=\frac{3}{(1-3x)^2}\]

OpenStudy (agreene):

that should be a one in the numerator for the derivative... it doesnt matter though, because any constant goes to 0.

OpenStudy (anonymous):

Now I'm just confused D:

OpenStudy (agreene):

Oh, I see where I messed up

OpenStudy (agreene):

\[\large\lim_{h\rightarrow0}\frac{\frac{1}{1-3(x+h)}-\frac{1}{1-3x}}{h}\] \[\large\lim_{h\rightarrow0}\frac{\frac{1}{1-3x-3h}-\frac{1}{1-3x}}{h}\] \[\large\lim_{h\rightarrow0}\frac{3}{1-3 h-6 x+9 h x+9 x^2}\] \[\large\frac{3}{\lim_{h\rightarrow0} (1-3 h-6 x+9 h x+9 x^2)}\] \[\frac{3}{1-6x+8x^2}\] \[\frac{3}{(1-3x)^2}\]